OpenStudy (anonymous):

The displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 4 sin πt + 3 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) (a) Find the average velocity during each time period? i) 1,2 ii) 1,1.1 iii)1,1.01 iv)1, 1.001????

OpenStudy (anonymous):

first we need \(s(1)=4\sin(\pi)+5\cos(\pi)=4\times 0+5\times -1=-5\) we keep that number then we need a calculator

OpenStudy (anonymous):

your next job is to compute \[\frac{s(2)-s(1)}{2-1}=\frac{s(2)+5}{1}=s(2)+5\] \[\frac{s(1.1)-s(1)}{1.1-1}=\frac{s(1.1)+5}{.1}=10\times (s(1.1)+5)\] and so on

OpenStudy (anonymous):

the first one you can do without a calculator, but not the others

OpenStudy (anonymous):

first one is \(10\) since \(\sin(2\pi)=0\) and \(\cos(2\pi)=1\) you get \(5+5=10\)

OpenStudy (anonymous):

second one needs help http://www.wolframalpha.com/input/?i=10*%284sin%281.1*pi%29%2B5cos%281.1*pi%29%2B5%29